OpenStudy (anonymous):

J is directly proportional to P. If J=80 when P=60, Find P when J=200

OpenStudy (anonymous):

J = kP where k is a constant 80 = 60k k = 4/3 J = 4P/3 200 = 4P/3 isolate, solve.